A student makes a $10.25 purchase at the bookstore with a $20 bill. The store has no bills and gives the change in quarters and fifty-cent pieces. there are 30 coins in all. how many of each kind are there?
Answer by checkley77(12844)

Q+H=30 OR Q=30-H
.25Q+.50H=20-10.25
.25(30-H)+.50H=9.75
7.5-.25H+.50H=9.75
.25H=9.75-7.50
.25H=2.25
H=2.25/.25
H=9 HALF DOLLARS WERE PART OF THE CHANGE.
Q=30-9=21 QUARTERS WERE PART OF THE CHANGE.
PROOF:
.25*21+.50*9=9.75
5.25+4.50=9.75
9.75=9.75